the jones family invests $11000 in two different money market accounts. one account pays 7% interest, and the other pays 8.5%. by the end of year the Jones family receives $860 in interest. How much did they invest in each account?
Answer by mananth(16949)

Part I 7.00% per annum ------------- Amount invested =x
Part II 8.50% per annum ------------ Amount invested = y
11000
Interest----- 860.00

Part I 7.00% per annum ---x
Part II 8.50% per annum ---y
Total investment
x + 1 y= 11000 -------------1
Interest on both investments
7.00% x + 8.50% y= 860
Multiply by 100
7 x + 8.5 y= 86000.00 --------2
Multiply (1) by -7
we get
-7 x -7 y= -77000.00
Add this to (2)
0 x 1.5 y= 9000
divide by 1.5
y = 6000
Part I 7.00% $ 5000
Part II 8.50% $ 6000

CHECK
5000 --------- 7.00% ------- 350.00
6000 ------------- 8.50% ------- 510.00
Total -------------------- 860.00

m.ananth@hotmail.ca